I'm not entirely sure how to do this problem. I need to set up the equation with one variable. I don't need to solve it.
A concert brings in $1148 in ticket sales. Regular tickets cost $8, while student tickets cost $5.
A total of 172 tickets were sold. How many regular tickets and student tickets were sold?

Answer by richwmiller(17219)

x+y=172,
x=172-y
y=172-x
8*x+5*y=1148
8*(172-y)+5*y=1148
solve for y
or
8*x+5*(172-x)=1148
solve for x